OpenStudy (anonymous):

what is an equation in which the answer is 28?

OpenStudy (2bornot2b):

\[x=28\] :)

OpenStudy (hoblos):

3x -2 = 2x + 26 is it like this??

OpenStudy (anonymous):

You mean a 3 step equation like the others?

OpenStudy (anonymous):

@2bornot2b: LOL

OpenStudy (2bornot2b):

You can't actually stop me from writing that equation

OpenStudy (anonymous):

Best equation ever 2bornot2b!!! Simple and does the job!!

OpenStudy (anonymous):

@2bornot2b: you really rock on that! :D

OpenStudy (anonymous):

I like x = 28

OpenStudy (anonymous):

Oh wait I have another one 28 = x :-)

OpenStudy (anonymous):

28 = x step 1) 28 * 2 = 56 = 2x step 2) 56 - 6 = 50 = 2x - 6 step 3) 50 / 2 = 25 = (2x - 6)/2 you can do that for any number

OpenStudy (anonymous):

give another one